width of a rectangle is 6 kilometers lesss than twice its length.if its area is 56 sqare kilometers
Answer by mathstutor494(120)

Let Length be "x"
Then width = 2x-6
Area = x*(2x-6) which is equal to 56 Sq Km
so the eqn is
x*(2x-6)=56
>2x^2 - 6x = 56
> 2x^2 - 6x - 56=0
> x^2 - 3x-28=0 (dividing by 2)
> x^2 -7x + 4x -28 = 0
> x(x-7) +4(x-7) = 0
> (x+4) (x-7) = 0
So x = -4, 7
As length can't be negative, so the answer is
Length =7km, Breadth = 8km